OpenStudy (anonymous):

8b^3/20b^3 asap help please!!!!!!!!

ganeshie8 (ganeshie8):

\(\large \frac{8b^3}{20b^3}\)

OpenStudy (anonymous):

b^3 is shared by the numerator and the denominator so they cancel each other out. It would be 8/20.

ganeshie8 (ganeshie8):

cancel b^3 top and bottom, also you can simplify the remaining :)

ganeshie8 (ganeshie8):

\(\large \frac{8b^3}{20b^3}\) \(\large \frac{8}{20}\)

ganeshie8 (ganeshie8):

i think 4 goes in both top and bottom. so.. ? let it go ?

ganeshie8 (ganeshie8):

\(\large \frac{8b^3}{20b^3}\) \(\large \frac{8}{20}\) \(\large \frac{2}{5}\)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

yeahhh thanks again!! i know you explained it to me the other day, but this one confuzzled the heck out of me. xD

ganeshie8 (ganeshie8):

hah takes time to fully digest i guess :) glad to hear u got it :))

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

so let it sink in.. and dont wry if u get few questions wrong... its okay !

ganeshie8 (ganeshie8):

just saying lol there are so many subjects to focus on... theres never enough time to i feel always :o

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

2b^6/5